OpenStudy (anonymous):

what are the equivalent fraction and decimal for 25%?

jimthompson5910 (jim_thompson5910):

x% = x/100 you take the number in front of the percent sign and put it over 100

OpenStudy (anonymous):

0.25 represents 25/100

jimthompson5910 (jim_thompson5910):

example: 98% = 98/100 reduce that to get 98/100 = 49/50 this is just an example and not the answer

OpenStudy (anonymous):

1/4